STUDY OF PHENYLACETYLENE BY CAVITY RING-DOWN SPECTROSCOPY

\underline{GARY V. LOPEZ}, PHILIP M. JOHNSON, TREVOR J. SEARS^{a} , Department of Chemistry, Stony Brook University, Stony Brook, New York 11794; and CHIH-HSUAN CHANG, Department of Chemistry, Brookhaven National Laboratory, Upton, New York 11973.

Cavity ring-down (CRD) measurements have been made on the S_{1}(Ã ^{1}B_{2}) ← S_{0}(X̃ ^{1}A_{1}) absorption spectrum of slit jet-cooled phenylacetylene (PA) in the 279 nm region as a further investigation of the photophysical properties of PA reported by Hosftein et al.^{b} The intensities in the new CRD data are compared with those obtained in the old REMPI experiments, and computational models.^{c} The results show that the strong bands in the REMPI spectrum are attenuated while the weak bands (a_{1} modes) remained the same when compared to CRD spectra. The comparison of the simulated spectrum to the CRD and REMPI spectra were satisfactory for most part when the a_{1} modes were scaled up by a factor of 3.24 and 3.55, respectively, giving evidence of some possible theoretical artifacts.^{c} The temperature of the slit jet-cooled PA was found to be in the range of 30 ± 5K by comparison with simulations of the rotational structure of the band origin of the S_{1} ← S_{0} transition. Additionally, many hot bands located near the band origin of the this transition have been assigned. Acknowledgments: Work at Brookhaven National Laboratory was carried out under Contract No. DE-AC02-98CH10886 with the U.S. Department of Energy and supported by its Office of Basic Energy Sciences, Division of Chemical Sciences, Geosciences and Biosciences.

aalso: Department of Chemistry, Brookhaven National Laboratory, Upton, New York 11973

bJ. Hofstein, H. Xu, T. J. Sears and P. M. Johnson, J. Phys. Chem. A 112, 1195 (2008)

cC. Chang , G. V. Lopez, T. J. Sears, and P. M. Johnson, J. Phys. Chem. A 114, 8262 (2010)